OpenStudy (anonymous):

An astronaut has a mass of 70kg. What is the weight of the astronaut on Earth at sea level?

OpenStudy (anonymous):

\(w=mg\), now what is \(g\) at sea level?

OpenStudy (anonymous):

w = 70 * 9.81

OpenStudy (anonymous):

weight = mass* acceleration due to gravity = 70*9.81=686.7 N